Question 6: The maintenance department of a large hospital uses about 1220 cases of liquid cleanser annually Ordering costs are $ 15, carrying costs are $ 25 per case a year, and the new price schedule is given in table below. Determine the optimal order quantity and the total cost. (15 points) Price Range 1 to 49 50 to 79 80 to 99 100 or more 31 S 34 32

I must be able to show my work.
Answer #1

Answer d: EOQ = 38.26 cases, Total cost= $ 2178.98

Explanation: S = Setup costs (per order) = $ 15
D = Demand (quantity sold per year) = 1220
H = Holding costs = $25

  • EOQ = √(2SD / H), = 38.26

Price of 1 case for ordering 38.26 cases= $34 (range 1 to 49)

Therefore, Total no. of orders= total demand/quantity=1226/38.26= 32.04

Therefore, total order cost= 32.04*$34= $1089.4

In EOQ, total ordering cost= total holding cost

  • Therefore, total cost= 2*total order cost= 2*1089.4= $ 2178.98

Please ask, if you have any doubts through comment section. Do rate the answer